OpenStudy (anonymous):

sin(x-180^0)tan405^0sin(90^0-x)/[sin^2(180^0+y)+cos^2(360^0-y)]cos(180^0-x)

OpenStudy (larseighner):

A or B or none of these? A. \[ \large {{\sin(x-180^\circ)\tan(405^\circ)\sin(90^\circ-x)} \over{\sin^2(180^\circ+y)}} +\cos^2(360^\circ-y)\cos(180^\circ - x) \] B. \[ \large {{\sin(x-180^\circ)\tan(405^\circ)\sin(90^\circ-x)} \over{\sin^2(180^\circ+y)}+\cos^2(360^\circ-y)\cos(180^\circ - x)} \]

OpenStudy (anonymous):

THXS